Solve the logarithmic equation and express irrational solutions in lowest radical form.
log (x+9) = log (6) - log (3x-27).
I appreciate any assistance with this!!

Answer by stanbon(48535)

log (x+9) = log (6) - log (3x-27).
log (x+9) = log[6/(3x-27)]
Take the anti-log of both sides to get:
x+9 = 6/(3x-27)
x+9 = 2/(x-9)
x^2-81=2
x^2-83=0
x= sqrt83 is the only solution
x=-sqrt83 results in taking the log of a negative number
which does not exist.
------------
Cheers,
Stan H.